The results of a nationwide aptitude test in mathematics are normally distributed with mean of 80 and standard deviation of 50.What is the percentile rank of a score of 84

1
Expert's answer
2022-04-06T16:26:12-0400

"P(X<84)=P(Z<\frac{84-80}{50})=P(Z<0.08)=0.5319."

Percentile rank of score 84 is 53.